# How To Embed Matters: Evaluation of EO Embedding Design Choices

### 11 Mar 2026 v1

@@ 5 Discussion and Conclusion
Our results show that compact EO embeddings can retain strong downstream utility, but their effectiveness depends on both encoder architecture and how information is preserved under spatial compression. Task-level analysis reveals varying difficulty across NeuCo-Bench tasks: semantic land-cover proportion targets are reliably captured, whereas continuous biophysical targets and atmospheric targets (Biomass, Clouds, Heat Island) are substantially more sensitive to architectural and representational choices. Differences in robustness, captured by the Q-score, can reveal variation even when mean $R^{2}$ appears saturated.
We employ the NeuCo-Bench framework, which relies on linear probing of 1D embeddings and limits downstream model complexity to isolate representation quality in a controlled and computationally efficient setting. Some regression targets, particularly continuous biophysical variables, contain inherent label uncertainty that may constrain absolute performance but affects all methods equally.
Backbone architecture and representation depth
play a decisive role. Consistent with prior findings on the global modeling capacity of transformer architectures [24, 23], ViTs outperform ResNets when using fixed-size embeddings for aggregated scene-level targets. However, optimal depth is task-dependent: ViT performance typically increases and saturates with depth, whereas ResNet features follow an inverted-U pattern, with intermediate stages substantially outperforming the final layer on continuous biophysical targets. This suggests that defaulting to final-layer embeddings can be suboptimal.
Self-supervised objectives and aggregation strategies
shape representation quality. Contrastive objectives (e.g., DINO) favor semantic signals, whereas reconstruction-based objectives (e.g., MAE, FGMAE) better capture continuous variation; no single objective dominates, and complementary strengths emerge. Mean pooling provides a robust aggregation strategy, while min and max pooling often discard informative structure. CLS pooling remains competitive for transformers, but concatenation shows limited complementarity beyond mean pooling alone.
Concatenation experiments
demonstrate that meaningful gains arise primarily from diversity across SSL objectives rather than token-aggregation strategies. When baseline task profiles differ, concatenation can consolidate complementary strengths and in some cases also yield per-task improvements beyond both baselines.
@@ Appendix A Final-Layer Embeddings: Full Per-Task Results
Overview. This section provides the complete per-task $R^{2}$ results for all evaluated GeoFM backbones (final layer embedding), self-supervised objectives, and pooling strategies: Table 1 reports mean pooling results for SSL4EO pretrained GeoFMs, similarly Table 2 reports max and Table 3 reports min pooling. Methods are sorted by average performance (Avg.) to make consistent cross-task trends explicit and to complement the Q-score–based radar plot shown in the main paper. Additionally, in Table 5 all pooling method results for TerraMind-Small are reported.
@@ Appendix C Per-Task Layerwise Results
Layer-wise analysis (ResNet). Figures 11 and 12 illustrate a more pronounced depth dependence for ResNet-50 models. While semantic and land-cover tasks benefit from deeper representations, multiple other tasks exhibit a clear performance drop at the final layer. Intermediate layers therefore remain are superior for several targets, supporting the main-paper observation that final-layer embeddings can reduce task-agnostic performance for convolutional backbones.
@@
Figure 11: Layer-wise per-task downstream performance ($R^{2}$) for ResNet-50 models pretrained on SSL4EO. Semantic and land-cover tasks show increasing and saturating trends similar to ViT models. In contrast, several other tasks exhibit a pronounced drop at the final layer, intermediate layers frequently remain competitive with the performance range of ViT final layer embeddings.

### 08 Apr 2026 v2

@@ 5 Discussion and Conclusion
Our results show that compact EO embeddings can retain strong downstream utility, but their effectiveness depends on both encoder architecture and how information is preserved under spatial compression. Task-level analysis reveals varying difficulty across NeuCo-Bench tasks: semantic land-cover proportion targets are reliably captured, whereas continuous biophysical and atmospheric targets (Biomass, Clouds, Heat Island) are substantially more sensitive to architectural and representational choices. Differences in robustness, captured by the Q-score, can reveal variation even when mean $R^{2}$ appears saturated.
We employ the NeuCo-Bench framework, which relies on linear probing of 1D embeddings and limits downstream model complexity to isolate representation quality. Some regression targets, particularly continuous biophysical variables, contain inherent label uncertainty that may constrain absolute performance but affects all methods equally.
Backbone architecture and representation depth play a decisive role. Consistent with prior findings on the global modeling capacity of transformer architectures [24, 23], ViTs outperform ResNets when using fixed-size embeddings for aggregated scene-level targets. However, optimal depth is task-dependent: ViT performance typically increases and saturates with depth, whereas ResNet features follow an inverted-U pattern, with intermediate stages substantially outperforming the final layer on continuous biophysical targets. This suggests that defaulting to final-layer embeddings can be suboptimal.
Self-supervised objectives and aggregation strategies shape representation quality. Contrastive objectives (e.g., DINO) favor semantic signals, whereas reconstruction-based objectives (e.g., MAE, FGMAE) better capture continuous variation; no single objective dominates, and complementary strengths emerge. Mean pooling provides a robust aggregation strategy, while min and max pooling often discard informative structure. CLS pooling remains competitive for transformers, but concatenation shows limited complementarity beyond mean pooling alone.
Concatenation experiments demonstrate that meaningful gains arise primarily from diversity across SSL objectives rather than token-aggregation strategies. When baseline task profiles differ, concatenation can consolidate complementary strengths and in some cases also yield per-task improvements beyond both baselines.
@@ Appendix A Final-Layer Embeddings: Full Per-Task Results
Overview. This section provides the complete per-task $R^{2}$ results for all evaluated GeoFM backbones (final-layer embedding), self-supervised objectives, and pooling strategies: Table 1 reports mean pooling results for SSL4EO pretrained GeoFMs, and similarly Table 2 and Table 3 report max and min pooling, respectively. Methods are sorted by average performance (Avg.) to make consistent cross-task trends explicit and to complement the Q-score–based radar plot shown in the main paper. Additionally, in Table 5 all pooling method results for TerraMind-Small are reported.
@@ Appendix C Per-Task Layerwise Results
Layer-wise analysis (ResNet). Figures 11 and 12 illustrate a more pronounced depth dependence for ResNet-50 models. While semantic and land-cover tasks benefit from deeper representations, multiple other tasks exhibit a clear performance drop at the final-layer. Intermediate layers therefore remain superior for several targets, supporting the main-paper observation that final-layer embeddings can reduce task-agnostic performance for convolutional backbones.
@@
Figure 11: Layer-wise per-task downstream performance ($R^{2}$) for ResNet-50 models pretrained on SSL4EO. Semantic and land-cover tasks show increasing and saturating trends similar to ViT models. In contrast, several other tasks exhibit a pronounced drop at the final-layer, intermediate layers frequently remain competitive with ViT final-layer embeddings.